### v1

# Predicting post-release defects with knowledge units (KUs) of programming languages: an empirical study

Abstract
Traditional code metrics (product and process metrics) have been widely used in defect prediction. However, these metrics have an inherent limitation: they do not reveal system traits that are tied to certain building blocks of a given programming language. We conjecture that taking these building blocks of a programming language into account can lead to further insights about a software system and improve defect prediction. Towards filling this gap, this paper reports an empirical study on the usage of knowledge units (KUs) of the Java programming language to predict post-release defects. A KU is a cohesive set of key capabilities that are offered by one or more building blocks of a given programming language. We operationalize our KUs via certification of the Java programming language. This study aims to understand whether we can obtain richer results in defect prediction (e.g., more accurate prediction of defect-prone code) when using KUs in combination with traditional code metrics. We select the defect dataset that was curated by (Yatish et al., 2019) which includes defect dataset for 28 releases of 8 Java software systems. We analyze the source code of these releases and extract traditional code metrics and the incidence of KUs. We find empirical evidence that KUs are different and complementary to traditional metrics, thus indeed offering a new lens through which software systems can be analyzed. Such a result motivates us to study the effectiveness of KUs in predicting defect-prone code. We build a defect prediction model called KUCLS, which leverages the KU-based features. Our KUCLS achieves a median AUC of 0.82 and significantly outperforms the CC_PROD (model built with product metrics). The normalized AUC improvement of the KUCLS over CC_PROD ranges from 5.1% to 28.9% across the studied releases. Although the KUCLS does not outperform the CC model (model built with product and process metrics), it maintains strong AUC values across the studied releases. Combining KUs with traditional code metrics leads to even higher-performing models (KUCLS+CC), significantly outperforming both the CC (an improvement ranges from 4.9% to 33.3%) and KUCLS models (an improvement ranges from 5.6% to 59.9%). Finally, we develop a cost-effective model (KUCLS_CC_COST_EFF) that significantly outperforms the CC. These encouraging results can be helpful to researchers who wish to further study the aspect of feature engineering and building models for defect prediction. Future work in this area thus (i) consider the KUCLS as a baseline model and (ii) include KU-based features in the design of models that predict post-release defects.
keywords:
Software engineering , Software metrics , code metrics , Java , certification exam , defect prediction , machine learning
[…]
[1] organization=Software Analysis and Intelligence Lab (SAIL), School of Computing, Queen’s University, city=Kingston, Ontario, country=Canada
[…]
[2] organization=York University, city=Toronto, Ontario, country=Canada
1 Introduction
Defect prediction has been a long-standing area of research in the field of software engineering, aimed at identifying potential defect-prone code in software systems (Koru et al., 2008; Nagappan et al., 2006; Moser et al., 2008; Rahman and Devanbu, 2013; Majumder et al., 2022). While the defect prediction field has seen substantial research in developing new machine learning approaches (e.g., bootstrapped by the recent advances and popularization of deep learning techniques (Hoang et al., 2019; Li et al., 2017; Chen et al., 2019; Zain et al., 2023)), the exploration of novel types of data for feature engineering remains less explored. The continual improvement in the defect prediction field demands innovation not just in modeling techniques but also in the data used for feature engineering. The more indicators researchers have, the more ammunition they have to understand defects and conjecture their real causes.
Traditional code metrics (e.g., product and process metrics) have been extensively used for defect prediction over several decades. These metrics characterize the size (e.g., LOC), complexity (e.g., cyclomatic complexity), and structure (e.g., the CK suite (Chidamber and Kemerer, 1994)) of object-oriented software systems. These metrics are programming-language agnostic, in the sense that their definitions are applicable to and computable for any object-oriented system (irrespective of the used programming language). As an inherent limitation, traditional code metrics do not reveal system traits that are tied to certain building blocks of a given programming language. As an illustrative example, let us consider a system written in Java and that makes use of the Java Concurrency API (i.e., a few classes from the java.util.concurrent package). Detecting the use of such an API reveals that concurrent code is a trait of that system. Such a trait is relevant: specific defects might arise because of it, not every developer is proficient in concurrent programming, and debugging a concurrent program is often more difficult than debugging a single-threaded one. Yet, traditional code metrics would not capture this concurrency trait.
Towards filling this gap, this paper reports an empirical study on the usage of knowledge units (KUs) of the Java programming language to predict post-release defects. As defined in our prior study (Ahasanuzzaman et al., 2024) a KU is a cohesive set of key capabilities that are offered by one or more building blocks of a given programming language (see Figure 1). The word “capabilities” refer to the main things that a programmer can do with a certain building block of a programming language. For instance, using the Concurrency API building block, a developer can create worker threads to execute tasks concurrently. Therefore, it is reasonable to assume that Java has a Concurrency KU, which includes a cohesive set of key concurrent program capabilities offered by the Concurrency API (building block). We operationalize our KUs using certification exams for the Java programming language.
In this paper, we conduct an exploratory study on the KUs of the Java programming language for predicting post-release defects. Our goal is to understand whether we can obtain richer results in defect prediction (e.g., more accurate prediction of defect-prone code) when using KUs in combination with traditional code metrics. We proceed as follows. First, we select the defect dataset that was curated by (Yatish et al., 2019). This dataset contains post-release defect data for 28 releases that span a total of 8 Java software systems. Then, we analyze the source code of each of the studied releases and collect the 28 KU-based features from each file of the defect dataset using our custom built KU detector. We also extract 59 traditional code metrics (54 product metrics and 5 process metrics) that have been used for defect prediction in prior studies. We start with a preliminary study to compare KU-based features with traditional code metrics. We apply the K-means algorithm to cluster the files of the studied releases using KUs. Subsequently, we recluster the systems using traditional code metrics and compare the two clusterings. We observe that KUs identify two times more cluster than traditional code metrics and the identified clusters of KUs do not overlap with the clusters of code metrics. Therefore, KUs are different from code metrics and can offer a new lens to analyze the source code of software systems. This study motivates us to study the effectiveness of KUs in post-release defect prediction. More specifically, we address the following six research questions (RQs):
RQ1: How effective are KU-based features for identifying defect prone code? First, we use the KU-based features to build a defect prediction model for each of the studied releases. We refer to this model as KUCLS. To compare the performance of the KUCLS, we build two baseline models using traditional code metrics: (i) CC_PROD model which is built with product metrics and (ii) CC model which is built with all the studied traditional code metrics (both process and product metrics). To compare the performance of KUCLS and any of the baseline model, we perform a statistical test using a two-sided Wilcoxon signed-rank test. Additionally, we employ Cliff’s delta (Long et al., 2003) as an effect size measure to assess the practical significance of the differences between the distributions. Our results indicate that:
KUCLS achieves a median AUC of 0.82. A defect prediction model with an AUC of 0.70 or higher is typically considered suitable for practical use (Jiarpakdee et al., 2020). The KUCLS outperforms the CC_PROD (median AUC of 0.80) in the defect prediction task. The normalized AUC improvement of the KUCLS over the CC_PROD ranges from 5.1% to 28.9% across the studied releases. Although the KUCLS does not outperform the CC model (median AUC of 0.84), it maintains strong AUC values across the studied releases.
RQ2: What are the important KU-based features of the KUCLS for predicting post-release defects? To determine the feature dimensions that significantly impact the KUCLS’s performance, we utilize SHAP (SHapley Additive exPlanation). SHAP is a robust, flexible, and widely used method for model interpretation (Molnar, 2020; Rajbahadur et al., 2022). We first compute the absolute SHAP values for each KU-based feature across all records in the dataset. Next, for each feature, we sum the SHAP values of all records in the dataset. This process is repeated for 100 bootstrapped datasets per release, resulting in a distribution of SHAP value for each feature. We then apply the Scott-Knott ESD (SK-ESD) technique (Ghotra et al., 2015; Mittas and Angelis, 2013; Tantithamthavorn et al., 2017) to rank features based on these distributions and obtain a feature ranking for each release. We take the ranking of a feature for all releases and produce a feature rank distribution for the feature. To summarize the results across all releases, we apply the Scott-Knott ESD technique again on the feature rank distributions, similarly to Ghotra et al. (2015). Our results indicate that:
Method Encapsulation, Inheritance, and Exception are identified as the top three most influential KU-based features of the KUCLS for predicting post-release defects consistently across all studied releases.
RQ3: Can KU-based features of the KUCLS provide insights into the incidence of post-release defects? To understand detailed insights into the factors influencing predictions for individual code segments, we perform an instance specific interpretation (a.k.a., local interpretations) using SHAP. We manually investigate a few buggy files, their corresponding issues in an ad-hoc manner and the SHAP values of the features of KUCLS and CC. Our results indicate that:
KU-based features offer insights into the characteristics of defect-prone files. Such insights are complementary to those offered by traditional code metrics.
RQ4: Can KUCLS be made more accurate by combining KU-based features with traditional code metrics? We combine the KU-baed features of the KUCLS and the features of the CC to determine whether we can obtain a higher-performing model. More specifically, we use all KU-based features of the KUCLS and all traditional code metrics features of the CC to build a combined model, which we call KUCLS+CC. We then apply the SK-ESD to compare the performance of KUCLS+CC with that of KUCLS and CC. To investigate the feature dimension of KUCLS+CC that most significantly influence predictions, we employ the same SHAP method from RQ2. Our results indicate that:
KUCLS+CC achieves a median AUC of 0.87 and outperforms both the KUCLS and CC across all the studied releases. The normalized AUC improvement of the KUCLS+CC over the CC ranges from 4.9% to 33.3% and the improvement over the KUCLS ranges from 5.6% to 59.9% across the studied releases. Although DDEV (a code metric feature) emerges as the most influential feature overall, Inheritance and Method Encapsulation KU-based features are the third ranked features indicating significant contribution of KU-based feature in predicting post-release defects.
RQ5: How different classification algorithms with hyper-parameter tuning impact the combined model for predicting post-release defects? We build the KUCLS+CC using a random forest classifier with default hyper-parameter values (scikit-learn python implementation). Since each classifier has its own strengths and weaknesses, the optimal performance of a prediction model can be achieved by choosing an appropriate classifier and tuning its hyper-parameters. To explore how different classifiers with hyper-parameter tuned parameters affect the KUCLS+CC model, we evaluate a mix of traditional and advanced classifiers: Naive Bayes (NB), k-nearest neighbor (KNN), Decision Tree (DT), Random Forest (RF), XGBoost (XGB), and LightGBM (LGBM). We build the KUCLS+CC models using each of these selected classifiers. To determine the most effective model hyper-parameter configuration, we utilize Scikit-learn’s GridSearchCV [^2] https://scikit-learn.org/stable/modules/generated/sklearn.model_selection.GridSearchCV.html. We compare these models and rank them using SK-ESD technique. Our results indicate that:
Among the studied classifiers, the KUCLS+CC that is built using a random forest classifier with default hyper-parameter values ranks as the highest performing model. Hence, we cannot further improve the performance of the KUCLS+CC from RQ4 by leveraging other classifiers and tuned hyper-parameters.
RQ6: How does the performance of a cost-effective combined model compare to that of the model built with traditional code metrics? A cost-effective model that uses fewer features yet maintains decent performance is thus valuable, as it reduces operational burdens associated with extensive feature engineering. Towards this goal, we build a cost-effective combined model, which we call KUCLS_CC_COST_EFF. This model is built using the top five features of the KUCLS and the top five code metric features of the CC. This results in a combined model with only 10 features, compared to the 59 features in CC and the 28 features in KULTC. Our results indicate that:
The KUCLS_CC_COST_EFF achieves a median AUC of 0.86 and outperforms the CC. The normalized AUC improvement of the KUCLS_CC_COST_EFF over the CC ranges from 5.5% to 33.3%. Therefore, we improve over the model built with traditional code metrics while still keeping feature engineering cost at a reasonable level.
Paper organization. Section 2 defines KUs and presents our approach for detecting them. Section 3 describes our data collection approach. Section 5 presents the motivation, approach, and findings that are associated with our research questions. Section 6 discusses the balance between performance improvement vs model complexity and directions to improvements of models. Section 7 describes the threats to the validity of our findings. Section 8 discusses related work. Finally, Section 9 concludes the paper.
2 Knowledge Units (KUs)
We introduce the concept of knowledge units (KUs) of programming language in our prior work (Ahasanuzzaman et al., 2024). To make this paper as self-contained as possible, in the following we provide a brief introduction to KUs (Section 2.1) and discuss our approaches for eliciting (Section 2.2) and detecting them (Section 2.3).
2.1 Definition
Every programming language has its own building blocks. The building blocks of a programming language are the “blocks” that developers use to “build” (write) code. In the case of Java, we consider them to be fundamental language constructs (e.g., arrays, if/else statements, try/catch statements) and APIs (e.g., Generic and Collection API, Concurrency API and String API). Each building block offers a set of capabilities essentially representing the actions a developer can perform using that building block. We define a Knowledge Unit (KU) as a cohesive set of key capabilities that are offered by one or more building blocks of a given programming language. The inclusion of “key” in our definition aims to ensure that KUs are centered around fundamental capabilities rather than those that are overly specific. Figure 1 illustrates key capabilities that are associated with the Concurrency API (left-hand side) and the synchronize language construct (right-hand side).
Figure 1: Our metamodel for knowledge units (KUs).
2.2 Eliciting KUs
Certification exams of a programming language (e.g., Oracle Java SE and Java EE certification exams for Java) aim to determine the proficiency of a developer in using the key capabilities offered by the building blocks of that language. Therefore, we can say that certification exams capture the KUs of a programming language. For instance, one of the topics covered in the Java SE 8 Programmer II certification exam is “Generics and Collections.” The subtopics of “Generics and Collections” that are covered in the exam include: (i) create and use a generic class, (ii) create and use ArrayList, TreeSet, TreeMap, and ArrayDeque objects, (iii) use java.util.Comparator and java.lang.Comparable interfaces, and (iv) iterate using forEach methods of a List. We interpret such a list of subtopics as the key capabilities that are offered by the Generics and Collections building block of the Java programming language. From this interpretation, we infer that Java has a Generics and Collections KU.
We reuse the KUs that we elicited in our prior work (Ahasanuzzaman et al., 2024). To elicit those KUs, we relied on the Oracle certification exams for the Java programming language. Oracle offers certification exams for different Java editions, such as Java Standard Edition (Java SE) and Java Enterprise Edition (Java EE). Specifically, for Java SE, there are two levels of certification exams:(i) Java SE 8 Programmer I Certification exam (Oracle, 2022a), and (ii) Java SE 8 Programmer II Certification exam(Oracle, 2022b). For Java EE, only one version of the exam is offered by Oracle: Oracle Certified Professional, Java EE Application Developer Certification exam (Oracle, 2022c). We elicited KUs directly from the topics outlined in these exams. In most cases, an exam topic was interpreted as a KU (e.g., the String Processing topic of Java SE 8 Programmer II Certification exam was interpreted as “Generics and Collections”) and its subtopics (e.g., create a generic class, use TreeSet, TreeMap and ArrayList and use java.util.Comparator) were interpreted as the key capabilities of that KU. At the end of this process, we elicited 28 KUs (Table 1). We refer to our prior work (Ahasanuzzaman et al., 2024) for the detailed description of this mapping process.
Table 1: The list of knowledge units (KUs) elicited for the Java programming language.
Knowledge unit (KU) Definition
[K1] Data Type The declaration and initialization of different types of variables (e.g., primitive type and parameterized type.)
[K2] Operator and Decision The usage of different Java operators (e.g., assignment, logical, and bit-wise operators) and conditional statements (e.g, if, if-else, and switch statements).
[K3] Array The declaration, instantiation, initialization and the usage of one-dimensional and multi-dimensional arrays.
[K4] Loop The execution of a set of instruction-s/methods repeatedly using for, while, and do-while statements and the skipping and stopping of a repetitive execution of instructions and methods using continue and break statements.
[K5] Method and Encapsulation The creation of methods with parameters, the use of overloaded methods and constructors, the usage of constructor chaining, the creation of methods with variable length arguments, and the usage of different access modifiers. This KU also describes encapsulation mechanisms, such as creating a set and get method for controlling data access, generating immutable classes, and updating object type parameters of a method.
[K6] Inheritance Developing code with child class and parent class relationship, using polymorphism (e.g., developing code with overridden methods), creating abstract classes and interfaces, and accessing methods and fields of the parent’s class.
[K7] Advanced Class Design Developing code that uses the final keyword, creating inner classes including static inner classes, local classes, nested classes, and anonymous inner classes, using enumerated types including methods and constructors in an enum type, and developing code using @override annotator.
[K8] Generics and Collection The creation and usage of generic classes, usage of different types of interfaces (e.g., List Interface, Deque Interface, Map Interface, and Set Interface), and comparison of objects using interfaces (e.g., java.util.Comparator, and java.lang.Comparable).
[K9] Functional Interface The development of code that uses different versions of defined functional interfaces (e.g., primitive, binary, and unary) and user-defined functional interfaces.
[K10] Stream API The development of code with lambda expressions and Stream APIs. This includes developing code to extract data from an object using peek() and map() methods, searching for data with search methods (e.g., findFirst) of the Stream classes, sorting a collection using Stream API, iterating code with foreach of Stream, and saving results to a collection using the collect method.
[K11] Exception The creation of try-catch blocks, the usage of multiple catch blocks, the usage of try-with-resources statements, the invocation of methods throwing an exception, and the use of assertion for testing invariants.
[K12] Date time API This KU refers to create and manage date-based and time-based events using Instant, Period, Duration, and TemporalUnit, and work with dates and times across timezones and manage changes resulting from daylight savings, including Format date and times values.
[K13] IO Reading and writing data from console and files and using basic Java input-output packages (e.g., java.io.package).
[K14] NIO Interacting files and directories with the new non-blocking input/output API (e.g., using the Path interface to operate on file and directory paths). Performing other file-related operations (e.g., read, delete, copy, move, and managing metadata of a file or directory).
[K15] String Processing The knowledge about searching, parsing, replacing strings using regular expressions, and using string formatting.
[K16] Concurrency The knowledge about functionalities that are related to thread execution and parallel programming. Some of these functionalities include: creating worker threads using Runnable and Callable classes, using an ExecutorService to concurrently execute tasks, using synchronized keyword and java.util.concurrent.atomic package to control the order of thread execution, using java.util.concurrent classes and using a parallel fork/join framework.
[K17] Database The creation of database connection, submitting queries and reading results using the core JDBC API.
[K18] Localization The knowledge about reading and setting the locale (Oracle defines a locale as “a specific geographical, political, or cultural region”) by using the Locale object, and building a resource bundle for each locale, and loading a resource bundle in an application.
[K19] Java Persistence The usage of object/relational mapping facilities for managing relational data in Java applications. With this knowledge, developers can learn how to map, store, update and retrieve data from relational databases to Java objects and vice versa.
[K20] Enterprise Java Bean The knowledge about managing server-side components that encapsulate the business logic of an application.
[K21] Java Message Service API The knowledge of how to create, send, receive and read messages using reliable, asynchronous, and loosely coupled communication.
[K22] SOAP Web Service Creating and using the Simple Object Access Protocol for sending and receiving requests and responses across the Internet using JAX-WS and JAXB APIs.
[K23] Servlet Handling HTTP requests, parameters, and cookies and how to process them on the server sites with appropriate responses.
[K24] Java REST API Creating web services and clients according to the Representational State Transfer architectural pattern using JAX-RS APIs.
[K25] Websocket Creating and handling bi-directional, full-duplex, and real-time communication between the server and the web browser.
[K26] Java Server Faces The knowledge about how to build UI component-based and event-oriented web interfaces using the standard JavaServer Faces (JSF) APIs.
[K27] Contexts and Dependency Injection (CDI) Managing the lifecycle of stateful components using domain-specific lifecycle contexts and type-safely inject components (services) into client objects.
[K28] Batch Processing Creating and managing long-running jobs on schedule or on demand for performing on bulk-data, and without manual intervention.
2.3 Detection of KUs
Analogously to our prior work (Ahasanuzzaman et al., 2024), we employ static analysis to detect KUs from a given Java project release. First, we parse all Java source files of the release using the Eclipse JDT framework (Eclipse, 2020). The JDT framework creates an Abstract Syntax Tree (AST) and provides visitors for every element of such tree, including names, types (class, interfaces), expressions, statements, and declarations. [^3] The grammar employed by JDT can be seen at https://github.com/eclipse-jdt/eclipse.jdt.core/blob/master/org.eclipse.jdt.core.compiler.batch/grammar/java.g
To detect a KU, we resort to the set of key capabilities of that unit (Tables 8). For instance, to detect the presence of the Generics and Collections KU, we use the visitors provided by JDT to detect the capabilities associated with such unit, namely: (i) create and use a generic class, (ii) create and use ArrayList, TreeSet, TreeMap, and ArrayDeque objects, and (iii) use java.util.Comparator and java.lang.Comparable interfaces. For instance, using type visitors, we can determine whether a given class is a generic class. Similarly, using statement visitors, we can find instantiations of an ArrayList. We also use JDT to collect binding information of methods, classes or variables (e.g., org.eclipse.jdt.core.dom.IVariableBinding resolves the binding information for a variable). In our study, we exclude type bindings that are related to third-party libraries since our primary interest lies in studying the KUs that are associated with the source code written by the developers of the studied projects.
We count the occurrences of KUs for every Java source file. For example, one of the capabilities associated with the Generics and Collections KU is to create a generic class. Hence, if we find a declaration of a generic class in a given Java file X.java belonging to a release R, we increment the counter for Generic and Collection KU that is associated with file X.java by one.
3 Data Collection
Our data collection process contains five main steps. First, we collect the defect dataset. Then, we retrieve the source code for the studied projects and releases. Next, we quantify the incidence of KUs in the source code of our studied projects. We also extract traditional code metrics from the source code of our studied releases. Finally, we create KU-based feature dataset and code metric dataset. Figure 2 presents an overview of our data collection process. We describe each step in more detail below.
Figure 2: An overview of our data collection process.
  • (Step 1) Collecting defect dataset.
    • (Step 1.1) Download the defect dataset. We choose the defect dataset that was curated by Yatish et al. (2019). Instead of relying on heuristic to identify post-release defects, the authors employ an approach that “leverages the earliest affected release that is realistically estimated by a software development team for a given defect”. We download the defect dataset from the article’s replication package [^4] https://github.com/awsm-research/replication-icse2019. The dataset contains post-release defect data for 28 releases that span a total of 8 Java project. Table 2 presents a statistical summary of the dataset.
    Table 2: A statistical summary of the dataset.
    Project
    Name
    Studied
    Releases
    Number of
    Java Files
    Line of
    Code
    Total Defects
    Reported
    Rate of Defects
    (across releases)
    ActiveMQ 5.0.0, 5.1.0, 5.2.0, 5.3.0 1,884-3,420 142-299K 3,157 6%-15%
    Derby 10.2.1.6, 10.3.1.4, 10.5.1.1 1,963-2,705 412-533K 3,731 14%-33%
    Groovy 1.5.7, 1.6.0.Beta_1, 1.6.0.Beta_2 755-884 74-90K 3,943 3%-8%
    HBase 0.94.0, 0.95.0, 0.95.2 1,059-1,834 246-534K 5,360 20%-26%
    Hive 0.9.0, 0.10.0, 0.12.0 1,416-2,662 287-563K 3,306 8%-19%
    JRuby 1.1, 1.4.0, 1.5.0, 1.7.0 731-1,614 105-238K 5,475 5%-18%
    Lucene 2.3.0, 2.9.0, 3.0.0, 3.1.0 8,052-2,806 101-242K 2,316 3%-24%
    Wicket 1.3.0.beta1, 1.3.0beta2, 1.5.3 1,672-2,578 109-165K 3,327 4%-7%
  • (Step 2) Collecting the source code data of the studied releases.
    • (Step 2.1) Download the source code of the studied releases. To download the source code of the studied software projects, we clone the repository of each project using the command git clone project_url where project_url is the url of the repository of the project in GitHub. Then, we extract the source code of a particular release using the command git checkout tags/tag_name. Here, tag_name is the name of one of the studied release for the project. The list of projects and their associated GitHub repository links can be found in our supplementary material.
  • (Step 3) Collecting KUs from the studied releases.
    • (Step 3.1) Determining the incidence of KUs in the source code. We count the occurrences of KUs in each Java file for every studied release in the defect dataset using our custom detector, which is built on top of Eclipse JDT (refer to Section 2.3).
  • (Step 4) Collecting code metrics from the studied releases.
    • (Step 4.1) Extract traditional code metrics from source code. In our study, we compare the incidence of KUs with traditional code metrics. To extract the code metrics we use the Understand tool [^5] https://scitools.com. We run the tool on the source code of each studied release and extract metrics for each Java file in the defect dataset. We extract 54 product metrics. Examples of product metrics: CountLineCode (number of lines of code), CountDeclMethodPublic (number of declared public methods in a class), and CountClassCoupled (number of other classes to which a class is coupled to). We also collect 5 process metrics: the number of commits (COMM), the number of lines added (ADDED_Lines), the number of lines deleted (DEL_LINES), the number of active developers (ADDEV) and the number of distinct developers (DDEV). Therefore, we extract a total of 59 metrics. Prior studies also use the subset of our extracted code metrics in their studies (Yatish et al., 2019; Nuñez-Varela et al., 2017; Jiarpakdee et al., 2020; Esteves et al., 2020). The full list of the collected 59 metrics for our study can be seen in B.
  • (Step 5) Creating the KU-based feature dataset and the code metrics dataset.
    • (Step 4.1) Vector representation of every file of the studied releases. Each file $F$ of a studied release $R$ is represented by two vectors in the dataset: one for KU-based features and one for code metrics. The vector for KU-based features has a length of 28, representing all the 28 identified KUs (See Table 8 in the Appendix). Each position in the vector stores the count value of a KU that is extracted from $F$. The code metrics vector has a length of 59, representing all the traditional code metrics that we study. Each position in the vector stores the value of a code metric that is extracted from $F$.
      The set of all KU vectors of all files of the studied releases can be seen as a table (dataset). We refer to it as the KU-based feature dataset ($D_{ku}$). Analogously, we refer to the set of all code metric vectors of all files of the studied releases as the code metrics dataset ($D_{cm}$). $Dataset(R,D_{ku})$ returns the KU-based feature dataset of a release $R$ and $Dataset(P,D_{cm})$ returns the code metrics dataset of the release $R$.
Summary Defect dataset: The defect dataset that was curated by Yatish et al. (2019). Data source for the source code: GitHub Pieces of data collected: Snapshots of the source code of the studied projects releases Datasets produced: KU-based feature dataset and traditional code metrics dataset
4 Preliminary Study: Do KUs offer a complementary perspective to the one offered by traditional code metrics?
Motivation. We compare KUs with traditional code metrics which are also derived from the same raw data (i.e., source code) of software systems. We conjecture that KUs provide a new perspective for the analysis of software code compared to traditional code metrics.
Approach. Our preliminary study is conducted in three steps, which are briefly explained below.
  • Step 1) Comparing KUs to code metrics via correlation analysis. As a first step to determine how KUs and code metrics differ from one another, we calculate the Spearman’s rank correlation coefficient ($rho$) between them. More formally, we calculate this correlation between all pairs of variables $x$ and $y$, where $x$ is a column in $D_{ku}$ and y is a column in $D_{cm}$. We choose the Spearman’s $\rho$ in lieu of other correlation coefficients (e.g., Pearson’s correlation) because the Spearman’s $\rho$ does not require normally distributed data. We assess Spearman’s $\rho$ using the following thresholds (Evans, 1996), which operate on 2 decimal places only: very weak for $|\rho|\leq 0.19$, weak for $0.20\leq|\rho|\leq 0.39$, moderate for $0.40\leq|\rho|\leq 0.59$, strong for $0.60\leq|\rho|\leq 0.79$, and large for $|\rho|\geq 0.80$.
  • Step 2) Clustering the files of the studied releases according to KUs and code metrics. We first cluster the files of a project using KUs. Next, we re-cluster the files of the project using code metrics and evaluate how the two clusterings differ. We use the k-means clustering algorithm (Comaniciu and Meer, 2002) to perform the clustering. K-means is one of the most popular clustering algorithms and has been widely used in software engineering research (Bishnu and Bhattacherjee, 2012; Yoon et al., 2007; Jureczko and Madeyski, 2010; Zhang et al., 2016). The detailed sub-steps are shown below:
    • Step 2a) Reducing the dimensions of the datasets. Many clustering algorithms (e.g., k-means algorithm) have inferior performance for high dimensional data  (Parsons et al., 2004). To overcome the dimensionality issue, we reduce the dimension of our data using PCA before running the k-means algorithm. We perform the PCA on both KUs dataset $(Dataset(P,D_{ku}))$ and code metric dataset $(Dataset(P,D_{cm}))$ of a project $P$ (see Section 3) using the PCA class of scikit-learn package [^6] https://scikit-learn.org/dev/modules/generated/sklearn.decomposition.PCA.html. We reduce the dimension of our data by selecting the first-n PCA components that account for at least 90% of the variance of the data.
    • Step 2b) Identify optimal number of clusters. The k-means algorithm takes the number of clusters as an input parameter (K). To identify the optimal number of clusters, we analyze the silhouette coefficient (Rousseeuw, 1987). The silhouette coefficient is a popular measure that has been widely used in prior studies (Nidheesh et al., 2020; Al-Subaihin et al., 2016; Panichella et al., 2013) for determining the optimal number of clusters. The silhouette coefficient ranges from -1.0 to 1.0. A coefficient close to one indicates that the object well-matches its own cluster and, simultaneously, poorly-matches its neighboring clusters. If most objects have a high coefficient, then the clustering configuration is deemed appropriate. We run k-means for K values ranging from 2 to 40 and calculate the mean silhouette for each K value. We determine the optimal number of clusters by selecting the K that maximizes the mean silhouette score.
    • Step 2c) Clustering the files of the studied releases. To cluster the files of the studied releases using the k-means algorithm, we employ the KMeans class of scikit-learn from python [^7] https://scikit-learn.org/1.5/modules/generated/sklearn.cluster.KMeans.html. We use the selected PCA components as the input dataset. Finally, we cluster the studied releases using the k-means algorithm and visualize the clusters using the fviz_cluster function of the factoextra R package (Kassambara and Mundt, 2017).
  • Step 3) Comparing the two clusterings. We compare the two clusterings produced in the prior step. This comparison is done both visually and then statistically through the computation of the adjusted Rand index. The Rand index refers to the probability that the two clusterings will agree on a randomly chosen pair of data points. The adjusted Rand index employs a correction-for-chance to the original Rand index. Intuitively, the adjusted Rand index is a measure of the similarity between two clusterings. The index ranges from zero (i.e., the two clusterings do not agree on any pair of points) to one (i.e., the clusterings are exactly the same).
Findings. Observation 1) The majority of KUs do not exhibit a strong correlation with code metrics. Table 3 presents the strength of correlation between KUs and code metrics. We observe that no KUs show a large correlation with code metrics. We also observe that only 5.6% of the identified KUs (e.g., Operator and Decision KU) show a strong a correlation with code metrics. However, the Spearman’s $\rho$ for all pairs consisting of a code metric and a KUs is below 0.60 for the majority of KUs (i.e., 94.4% of the identified KUs). Therefore, the majority of KUs do not exhibit a strong correlation with code metrics. This result indicates that KUs can provide an additional perspective for the analysis of software code.
Table 3: The strength of correlation between KUs and code metrics.
Strength of correlation The percentage of KUs
Large correlation ($|\rho|\geq 0.80$) 0%
Strong correlation ($0.60\leq|\rho|\leq 0.79$) 5.6%
Moderate correlation ($0.40\leq|\rho|\leq 0.59$) 22.2%
Weak correlation ($0.20\leq|\rho|\leq 0.39$) 33.3%
Very weak correlation ($|\rho|\leq 0.19$) 38.9%
[…]
Figure 3: Identified clusters for knowledge units (KUs) and code metrics of the “Wicket-1.3.0-beta2” release.
Observation 2) KUs identify two times more clusters than code metrics. We observe that the clustering based on KUs produce 4 (median) clusters, whereas the clustering of code metrics produce e (median) clusters. In other words, the Java files of the studied releases can be seen as belonging to two times more clusters through the lens of KUs than the clusters that are identified by code metrics. Figure 3 depicts the identified clusters for KUs and code metrics for the “Wicket-1.3.0-beta2” [^8] https://github.com/structurizr/java. We observe that the clustering based on the lens of KUs produce 4 clusters and code metrics produce 2 clusters.
Finally, we observe that the median value of adjusted Rand index between the two clusterings across the studied releases is 0.54 which is lower than 1. In other words, the probability that the two clusterings will agree on a randomly chosen studied release is 54% (after correction-for-chance adjustment). Since a Rand index value other than 1 indicates differences in clusters, we conclude that the clusters of KUs are not the same as those of code metrics.
Figure 4: Visualization of clusters for KUs and code metrics. Each polygon represents a cluster of KUs and the dotted points of colors represent files in the clusters of code metrics.
Observation 3) The identified clusters via code metrics do not completely overlap with the identified clusters via KUs. Figure 4 shows the overlap between the two clusterings of the “Wicket-1.3.0-beta2” project. Each polygon represents a cluster of KU and each colored point represents a file in one of the clusters of the code metric clustering. As shown in Figure 4, cluster 2, 3 and 4 of KUs (i.e., the polygons) contain files of different clusters of code metrics (i.e., each polygon contains different coloured points). For example, cluster 4 of KUs contains 33.3% files of cluster 0 of code metrics and 67.7% files of cluster 1 of code metrics. Therefore, these KU clusters are not completely overlapped by a single cluster of code metrics (i.e., a KU cluster contains files of more than one cluster of code metrics). This result is also an indication that KUs offer a complementary perspective for analyzing software code.
Figure 5: The distribution of the percentage of KU clusters that are not overlapped by a single cluster of code metrics.
Figure 5 presents the distribution of the percentage of KU clusters that are not overlapped by a single cluster of code metrics across the studied releases. The median of the percentage of KU clusters that are not overlapped by a single code metric cluster is 81.67%, indicating that code clusters do not completely overlap with the KU clusters (i.e., polygons have a somewhat mix of different colored dots). Such a result indicates that there exists KU clusters that are not uniquely described by a single cluster of code metrics. This result is also an indication that KU clusters can capture a different perspective of software systems from the identified clusters via code metrics.
Summary Preliminary Study: Do KUs offer a complementary perspective to the one offered by traditional code metrics? KUs offer a new lens through which the code of software systems can be analyzed, revealing a perspective that is complementary to that yielded by traditional code metrics. In particular: None of the KUs exhibit a large correlation with code metrics. Through the lens of KUs, the files of the studied projects could be grouped into two times more cluster than the identified clusters via code metrics. The identified clusters using code metrics do not completely overlap with the identified clusters using KUs.
5 Findings
In this section, we address our research questions. For each question, we discuss our motivation for studying it, the approach that used to answer it, and the findings that we observe.
5.1 RQ1: How effective are KU-based features for identifying defect prone code?
5.1.1 Motivation
From our preliminary study, we conclude that KUs offer a new lens through which the code of software systems can be analyzed. This conclusion motivates us to study the effectiveness of KUs in post-release defect prediction problem. Traditionally, defect prediction has relied heavily on code metrics, such as process and product metrics (Gyimothy et al., 2005; Moser et al., 2008; Zhang, 2009; Shihab et al., 2010). We believe that exploring novel types of data for feature engineering is as important as developing or assessing new machine learning approaches for defect prediction. The more indicators researchers have, the more ammunition they have to understand bugs and conjecture their real causes.
5.1.2 Approach
First, we build a defect prediction model using the KU-based features for each of the studied releases. We refer to this model as KUCLS. Next, we evaluate the performance of the KUCLS. Finally, we compare the performance of the KUCLS to those models that are built using traditional code metrics. In the following, we explain each of the steps that we follow during this process.
Figure 6: An overview of building the KUCLS.
  • (Step 1) Construct the KUCLS model. In this step, we construct and evaluate KUCLS prediction model (for clean/buggy classes). An overview of our approach for building the KUCLS can be seen in Figure 6.
    • (Step 1.1) Select non-correlated features. Prior to building any defect prediction model, it is recommended to mitigate correlated features (Jiarpakdee et al., 2019) to better interpret the results of a model. Therefore, we want to check the performance of our studied models using non-correlated features. We select non-correlated features using the approach proposed by Jiarpakdee et al. (2018). This approach is provided by the AutoSpearman function in the python implementation [^9] https://xai4se.github.io/defect-prediction/data-preprocessing.html. AutoSpearman automatically selects non-correlated metrics based on two analyses: (1) a Spearman rank correlation test and (2) a Variance Inflation Factor analysis. We apply the AutoSpearman on KU-based feature dataset, which resulted in 28 non-correlated KU-based features (i.e., no KU-based feature is removed).
    • (Step 1.2) Generate bootstrap samples. To estimate the performance of our models in practice, we employ the out-of-sample bootstrap validation technique with 100 repetitions (Tantithamthavorn et al., 2017; Efron, 1983). In this technique, we randomly sample dataset with replacement to create 100 bootstrap samples. Each bootstrap sample have the same number of data points as the original dataset, but some points may appear multiple times while others may be excluded. These excluded data points from each bootstrap sample form the corresponding out-of-sample data, which we use for testing our models. We train our models on each bootstrap sample and test its performance on the associated out-of-sample data.
    • (Step 1.3) Construct the KUCLS model. We use the non-correlated KU-based features to construct our KUCLS model. In terms of the machine learning algorithm, we choose a random forest classifier with default parameters values (scikit-learn python implementation). This parameter setting of random forest classifier is identified as the optimal choice for achieving the best performance in classifying post-release defects (Yatish et al., 2019).
    • (Step 1.4) Evaluate the KUCLS model. To evaluate the accuracy of our defect prediction models, we calculate the Area Under the receiver operator characteristic Curve (AUC) as it is a robust, threshold-independent measure (Tantithamthavorn and Hassan, 2018). AUC value ranges from 0 to 1, with 0.5 being as good as random guessing. The AUC values closer to one indicate a model that is good at distinguishing the class of interest (buggy) from the other class (clean). Values closer to zero indicate a model that is good at predicting the class that is not of interest (clean). Since we are employing an out-of-sample bootstrap model validation with 100 repetitions, we obtain 100 AUC values for each prediction model that we evaluate.
  • (Step 2) Construct baseline models. Our main goal in RQ1 is to understand the efficacy of KUs in predicting software defects. Ultimately, we wish to understand whether we can obtain richer results (e.g., more accurate predictions) when using KUs in lieu of baselines. To keep things fair and more straightforward, we compare features that are derived from the same raw data (i.e., source code). Since KUs are derived from the source code, we compare KUs with traditional code metrics which are also derived from the source code. Moreover, previous studies widely used these subsets of our extracted code metrics for building and analyzing defect prediction models (Zhang et al., 2016; Yatish et al., 2019; Nuñez-Varela et al., 2017; Jiarpakdee et al., 2020; Laaber et al., 2021; Esteves et al., 2020). Therefore, we use the traditional code metrics and construct two baseline models: (i) a model built with product code metrics (CC_PROD), and (ii) a model built with both product and process metrics (CC). These baseline models are constructed following the procedure described in Step 1.
  • (Step-3) Compare models’ performance. To check whether the performance difference between a given pair of models (e.g., KUCLS vs CC_PROD) is statistically significant, we apply the Wilcoxon signed-rank test. The Wilcoxon signed-rank test is a non-parametric statistical test for paired data. We infer that there is a statistically significant difference between the two input distributions when the p-value computed by the test is smaller than 0.05 (i.e., $alpha=0.05$). In addition, we also use the Cliff’s delta ($d$) effect size measure to quantify the practical difference between the two distributions (Long et al., 2003). We use the following thresholds for interpreting $d$ (Romano et al., 2006): negligible for $|\delta|\leq 0.147$, small for $0.147<|\delta|\leq 0.33$, medium for $0.33<|\delta|\leq 0.474$, and large otherwise.
    To evaluate a model’s performance improvement compared to a baseline, we calculate the normalized AUC improvement metric. This metric measures the extent of improvement achieved by the model relative to the maximum possible improvement. It is calculated using the following formula:
    $\scriptstyle Normalized\ AUC\ improvement=\frac{AUC\ of\ ProposedModel-AUC\ of\ BaselineModel}{1-AUC\ of\ BaselineModel}\times 100\%$
    Here, 1 represents the highest possible AUC score. For example, to calculate the normalized AUC improvement of KUCLS over CC_PROD, where the AUC of KUCLS is 0.81 and the AUC of CC_PROD is 0.75, the calculation is as follows: the normalized AUC improvement is (0.81-0.75) / (1 - 0.75) $\times$ 100 % = 24%. This result indicates that KUCLS’s performance is 24% closer to the maximum possible improvement compared to the baseline. Since we construct 100 bootstrap models, we compute the average normalized AUC improvement across these models to provide a comprehensive measure of overall performance improvement.
Table 4: P-value, effect size, and median AUC value of KUCLS, CC_PROD, and CC defect prediction models.
Statistical Test
Median AUC KUCLS vs CC_PROD KUCLS vs CC
Release KUCLS CC_PROD CC P-value Effect Size P-value Effect Size
activemq-5.0.0 0.82 0.81 0.89 <0.05 M(0.44) <0.05 L(-1.00)
activemq-5.1.0 0.83 0.80 0.82 <0.05 L(0.59) <0.05 N(0.12)
activemq-5.2.0 0.83 0.80 0.85 <0.05 M(0.46) <0.05 L(-0.59)
activemq-5.3.0 0.82 0.79 0.83 <0.05 M(0.40) <0.05 M(-0.40)
activemq-5.8.0 0.85 0.84 0.86 <0.05 S(0.26) <0.05 M(-0.44)
derby-10.2.1.6 0.74 0.77 0.85 <0.05 L(-0.87) <0.05 L(-1.00)
derby-10.3.1.4 0.84 0.85 0.85 <0.05 S(-0.32) <0.05 L(-0.49)
derby-10.5.1.1 0.82 0.82 0.84 >0.05 N(0.01) <0.05 L(-0.62)
groovy-1.5.7 0.82 0.78 0.79 <0.05 S(0.29) <0.05 S(0.20)
groovy-1.6.BETA_1 0.82 0.81 0.87 <0.05 S(0.18) <0.05 L(-0.60)
groovy-1.6.BETA_2 0.85 0.83 0.91 <0.05 L(0.61) <0.05 L(-0.77)
hbase-0.94.0 0.72 0.78 0.84 <0.05 L(-0.95) <0.05 L(-1.00)
hbase-0.95.0 0.72 0.76 0.81 <0.05 L(-0.93) <0.05 L(-1.00)
hbase-0.95.2 0.71 0.69 0.72 <0.05 S(0.19) <0.05 L(-0.87)
hive-0.10.0 0.72 0.78 0.83 <0.05 L(-0.96) <0.05 L(-1.00)
hive-0.12.0 0.80 0.72 0.84 <0.05 L(0.99) <0.05 L(-1.00)
hive-0.9.0 0.79 0.82 0.90 <0.05 L(-0.96) <0.05 L(-1.00)
jruby-1.1 0.86 0.82 0.85 <0.05 L(0.54) <0.05 S(0.15)
jruby-1.4.0 0.83 0.81 0.84 <0.05 S(0.17) <0.05 N(-0.13)
jruby-1.5.0 0.84 0.83 0.86 <0.05 S(0.27) <0.05 S(-0.29)
jruby-1.7.0.preview1 0.84 0.81 0.83 <0.05 L(0.56) <0.05 S(0.17)
lucene-2.3.0 0.85 0.82 0.91 <0.05 L(0.65) <0.05 L(-1.00)
lucene-2.9.0 0.79 0.78 0.82 <0.05 S(0.19) <0.05 L(-0.8)
lucene-3.0.0 0.91 0.89 0.9 <0.05 L(0.60) <0.05 S(0.33)
lucene-3.1.0 0.74 0.72 0.76 <0.05 M(0.39) <0.05 L(-0.71)
wicket-1.3.0-beta2 0.81 0.80 0.82 <0.05 S(0.21) <0.05 S(-0.26)
wicket-1.3.0-incubating-beta-1 0.83 0.84 0.87 >0.05 N(-0.07) <0.05 L(-0.61)
wicket-1.5.3 0.81 0.77 0.79 <0.05 L(0.61) <0.05 M(0.43)
Median 0.82 0.80 0.84
[…]
5.1.3 Findings
Observation 4) The minimum AUC of the KUCLS is above 0.70 (i.e., AUC > 0.70), indicating that KUs are a good candidate to preeict post-release defects. Table 4 shows the P-value, effect size, and median AUC value of the prediction models for each studied release. The KUCLS column refers to the AUC of the models built with KU-based features. We observe that the AUCs of the KUCLS are always above 0.70 (minimum is 0.72). In particular, the median AUC of the KUCLS across all the studied releases is 0.82. A defect prediction model is typically considered for use in practice when its AUC is higher than or equal to 0.70 (Jiarpakdee et al., 2020). This result highlights that KUs are a good candidate to classify post-release defects.
Observation 5) The KUCLS outperforms the CC_PROD in classifying post-release defects. As we observe in Table 4, the median of the CC_PROD column is 0.80, whereas the median of the KUCLS column is 0.82. We also compare the individual releases, and we observe that the KUCLS outperforms the CC_PROD in the majority of the studied releases. We observe that the KUCLS outperforms the CC_PROD in 20 of the 28 releases (see green rows in Table 4 of the column named “KUCLS vs CC_PROD”), the majority of them with a large and medium effect size. The effect size is negligible in 2 out of 28 releases, indicating that both models perform the same in those cases. However, for the remaining 6 releases, the KUCLS does not outperform the CC_PROD, but maintain a good AUC value (the median AUC of KU models for these 6 releases is 0.73). In summary, the KUCLS wins in 20 releases, the CC_PROD wins in 6 releases, and a draw occurs in the remaining cases. We also calculate the normalized AUC improvement between the KUCLS and the CC_PROD. The normalized AUC improvement of the KUCLS over the CC_PROD is between 5.1% and 28.9% across the studied releases. These results show that the KUCLS outperform the CC_PROD.
Observation 6) The KUCLS generally does not outperform the CC in most of the studied releases. However, the KUCLS still achieves strong AUC values. As we observe in Table 4 the median of the CC column is 0.84, whereas the median of the KUCLS column is 0.82. We observe that the KUCLS outperforms the CC in 5 of the 28 studied releases. For the remaining releases (23 of the 28 studied releases), CC models outperform KUCLS models. The median AUC values of KUCLS for those 23 releases is 0.82. Therefore, while the KUCLS does not outperform the CC, the KUCLS still achieves strong AUC values.
Summary RQ1: How effective are KU-based features for identifying defect prone code? KUs are good candidates for predicting bug-prone code. In particular: The KUCLS can classify post-release defects with a minimum AUC of 0.73 and a median AUC of 0.82. The KUCLS outperforms the CC_PROD in the defect prediction task. The normalized AUC improvement of the KUCLS over the CC_PROD ranges from 5.1% to 28.9% across the studied releases. While the KUCLS does not outperform the CC, it maintains strong AUC values.
5.2 RQ2: What are the important KU-based features of the KUCLS for predicting post-release defects?
5.2.1 Motivation
Studying the feature importance of the KUCLS is essential to identify which KU-based feature contribute most significantly to predicting post-release defects. This analysis not only enhances our understanding of the model’s decision-making process but also highlights the practical relevance of specific KUs. From an operational perspective, identifying key features streamline the feature engineering process and improves the model’s practicality for deployment in real-world scenarios.
5.2.2 Approach
To determine the KU-based features that have the most significant impact on predictions, we conduct a model interpretation analysis utilizing SHAP (SHapley Additive exPlanation). SHAP, a widely adopted method for model interpretation, leverages game theory to provide an optimal framework for quantifying each feature’s contribution to the model’s predictions (Molnar, 2020). SHAP values can be positive or negative, signifying the direction of the feature’s influence on the prediction. A positive SHAP value indicates that the feature increases the likelihood of the predicted outcome (e.g., classifying a file as buggy in a defect prediction model). In contrast, a negative SHAP value means that the feature reduces the likelihood of the predicted outcome or increases the likelihood of the opposite class (i.e., not buggy class). The strength of a feature is determined by the magnitude of SHAP values. Specifically, the absolute SHAP value of a feature indicates its influence on the prediction, with higher values signifying greater importance. We use the SHAP package of python to perform the SHAP analysis (CRAN, 2020).
We compute KU-based feature importance for the KUCLS as follows. First, we apply SHAP to the KUCLS and calculate the absolute SHAP value for each feature across every file (i.e., record) in the dataset of a release. Next, for each feature, we sum the SHAP values of all files (i.e., records) in the dataset. We follow this approach for all the 100 bootstraps of every release. As a result we obtain a distribution of the sum of SHAP values for every feature across all releases. Then, we apply the Scott-Knott ESD technique on the distribution of the sum of SHAP values generated from the 100 bootstraps of every release. As a result, we obtain a feature ranking for each release. For example, in release jruby-1.1, the most important feature of the KUCLS is Inheritance (rank 1), the second most important feature is Exception (rank 2) and so on. There can be ties, i.e., two features with same importance are given the same rank. By taking all releases, we can produce a feature rank distribution for each feature. To summarize the results across all releases, we apply the Scott-Knott ESD technique again on the feature rank distributions, similarly to Ghotra et al. (2015). The SK-ESD is a method of multiple comparison that leverages a hierarchical clustering to partition the set of treatment values (e.g., medians or means) into statistically distinct groups with non-negligible difference (Mittas and Angelis, 2013). As a result, we obtain a final rank for each feature. We use Tim Menzies’s implementation of the Scott-Knott ESD technique, which employs non-parametric statistical tests and the Cliff’s Delta measure of effect size (Menzies, 2020).
Figure 7: The rank distribution of features of the KUCLS for defect classification. The number inside the square brackets indicates the final rank of the feature after applying Scott-Knott ESD for the second time.
5.2.3 Findings
Observation 7) Method Encapsulation is the most important KU-based feature across all studied releases of the KUCLS for predicting post-release defects. Figure 7 shows the rank distribution of each feature (final rank is shown in brackets). We observe that Method Encapsulation is the top-ranked KU in the KUCLS. Indeed, the boxplot reveals that Method Encapsulation KU is the most important feature in most of our studied releases as the median value is one and the interquartile range is narrow (most ranks are one). Inheritance KU is ranked second and Exception KU is the third-ranked feature. We then have two KU-based features ranked forth, namely: Advanced Class Design and LOOP. Of all these features, we note that Advanced Class Design KU is the one with the most stable rank (smallest interquartile range). The other KU is a little less stable, meaning that its rank varies more across the studied releases. Some advanced KUs derived from Java SE II certification exams also rank among the top ten most important features. For instance, Generics and Collections and Java IO are both ranked eighth and exhibit stable rankings (small interquartile range). While Concurrency KU ranks eleventh, its ranking shows greater variability with a large interquartile range. KUs related to building enterprise-level applications in Java, such as Enterprise JavaBeans and Java Message Service (JMS), which are derived from Java EE certification exams, are generally ranked lower in importance, with most being ranked 14th. However, they also have significant contributions in predicting post-release defects. These advanced Java EE features are less frequently used in the studied systems compared to more common KU-based features like Data Type and Loop. As a result, they may have minimal influence on the global defect prediction model.
Summary RQ2: What are the important KU-based features of the KUCLS for predicting post-release defects? Method Encapsulation, Inheritance, and Exception are identified as the top three most influential KU-based features of the KUCLS for predicting post-release defects consistently across all studied releases.
5.3 RQ3: Can KU-based features of the KUCLS provide insights into the incidence of post-release defects?
5.3.1 Motivation
While global interpretations of defect prediction models provide a broad overview of the relationships between features and defects, they often fail to capture the nuances of specific instances. In contrast, instance-specific interpretations (a.k.a., local interpretations) offer detailed insights into the factors influencing predictions for individual source code (i.e., file). Such insights can reveal unique combinations of traits and contexts that contribute to defects, enabling a more targeted understanding of their occurrence. Hence, we investigate whether an instance-specific interpretation from the KUCLS can provide insights into the occurrence of defects or the context in which defects happen. We also compare this instance-specific interpretation with that given by the CC model. Our conjecture is that the two explanations can complement each other.
5.3.2 Approach
We select the activemq-5.8.0 release to study this research question. Our justification is that the median performance of the KUCLS and the CC are both above 0.85 for this release. To pinpoint and contrast the explanations given by each model, we perform a local interpretation analysis using SHAP. With SHAP, one can perform instance-specific (i.e., file level) interpretations. To select the instance, we first investigate the issues that are linked to the files. We parsed the commit messages of all commits that were submitted during the development of the activemq-5.8.0 release and extracted issues IDs from these messages. As a result, we established a mapping between defect-prone files in activemq-5.8.0 (as per our defect dataset) and issue IDs. We manually investigate a few defect-prone files and their corresponding issue tracking IDs in an ad-hoc manner and chose one case to showcase its results in this RQ. We choose the case where the priority of the issue was Major and the description of both the issue and the fix were easily understandable and clearly identifiable. Our chosen file was ActiveMQSession.java and the associated issue ID was AMQ-4634101010https://issues.apache.org/jira/browse/AMQ-4634. According to the issue reporter, the issue seems to be related to how ActiveMQ handles transactions in a Java Message Service (JMS) environment when working with distributed transactions.
Figure 8: The SHAP values of the top ten features of the CC and KUCLS models for their contribution to classifying the ActiveMQSession.java file as defect-prone.
5.3.3 Findings
Observation 8) KU-based features of the KUCLS offer insights into the characteristics of defect-prone files, which can complement the insights given by the CC model. Figure 8 presents the result of the SHAP analysis of the defective file for the CC and KUCLS models. As shown in Figure 8(a), the top-3 features that more strongly drive the CC model’s classification decision are: ADEV (i.e., the number of active developers), CountDeclMethodPublic (i.e., the number of local public methods) and CountClassCoupled (i.e., the number of other classes coupled to). The positive SHAP values of these code metrics highlight the positive contributions of the file to be classified as defect. Notably, the first two metrics, ADEV and CountDeclMethodPublic, have significantly higher positive SHAP values (both exceeding 0.10), compared to CountClassCoupled, which has a SHAP value of only 0.03. We do observe a high number of active developers (30), public methods (28) and coupled classes (32) in this class.
Figure 8(b) shows the KU-based features that drive the KUCLS model’s decision. The top three KU-based features that positively contribute to classifying the file as defect-prone are: Message Service, String Processing, and Java IO. Among these, Message Service has the highest positive SHAP value (0.14), significantly higher than the SHAP values of the other KU-based features. This result indicates that the Message Service KU is strongly associated with classifying this file as defect-prone. By manually investigating the patch [^11] https://shorturl.at/Kkc3L, we note that one of the changed methods (run(), line 860 does use Message Service APIs (an instantiation of an object of MessageAck class). In addition, the defect-prone class file instantiates and uses a set of methods and subclasses of several Message Service APIs (e.g., Session, Message, QueueSender and QueueReceiver). Therefore, it becomes clear that the defect-prone class (and the defect-prone method) do operate in a context where the concept of message service matters. Hence, we conclude that, in this case, KU-based features could provide further insights into the context in which the defect happened.
By combining the interpretations from code metrics and KU-based features, we now know two things: (i) the defect-prone class consists of too many methods and dependencies (ii) the defect-prone class and method operate in a context where message service matters. While these explanations might not necessarily lead to the discovery of causal relationships (which is a limitation that is actually inherent to the interpretation of machine learning models), we believe that they tell more about the context in which the defect happened and can thus inform future maintenance and refactoring decisions surrounding a given defect-prone class. We highlight that the instance discussed in this section is just one illustrative example of a concrete scenario. Nevertheless, the same approach (SHAP) could be used to investigate any defect-prone file from our dataset. Having said that, we encourage future work to investigate more defects and ask for feedback from the developers who worked on resolving these defects.
Summary RQ3: Can KU-based features of the KUCLS provide insights into the incidence of post-release defects? Yes. KU-based features offer insights into the characteristics of defect-prone files. Such insights are complementary to those offered by traditional code metrics.
5.4 RQ4: Can KUCLS be made more accurate by combining KU-based features with traditional code metrics?
5.4.1 Motivation
This research question investigates whether the predictive performance of the KUCLS for post-release bugs can be improved by integrating them with traditional code metrics. The underlying motivation is based on the hypothesis that combining diverse sets of features – KUs and traditional code metrics – can provide a more comprehensive understanding of the factors influencing post-release bugs. Such integration may enable the combined model to capture a wider range of predictive indicators, thereby enhancing its accuracy in identifying potential bugs.
5.4.2 Approach
We build a combined model named KUCLS+CC for predicting post-release bugs using all studied features of the KUCLS and all studied traditional code metrics of the CC (process and product metrics). Following the outlined steps in Step 1 of RQ1, we build the KUCLS+CC models for all studied releases. We compare the performance of the KUCLS+CC model with the performance of the KUCLS and CC models. We rank the studied prediction models (KUCLS, CC_PROD, CC and KUCLS+CC) using SK-ESD (Ghotra et al., 2015; Mittas and Angelis, 2013; Tantithamthavorn et al., 2017). We also analyze the feature importance of the KUCLS+CC using the same SHAP analysis method explained in Section 5.2.
5.4.3 Findings
Figure 9: The distribution of AUC of KUCLS, CC, CC_PROD and the combined KUCLS+CC models. The models are grouped based on their performance rankings determined by the Scott-Knott ESD (SK-ESD) method, where a lower SK-ESD rank indicates a better-performing model.
Table 5: P-value, effect size, and median AUC value of KUCLS, CC, and KUCLS+CC defect classification models.
Statistical Test
Median AUC KUCLS+CC vs KUCLS KUCLS+CC vs CC
Release KUCLS+CC KUCLS CC P-value Effect Size P-value Effect Size
activemq-5.0.0 0.91 0.82 0.89 <0.05 L(1.0) <0.05 L(0.82)
activemq-5.1.0 0.85 0.83 0.82 <0.05 M(0.46) <0.05 L(0.56)
activemq-5.2.0 0.87 0.83 0.85 <0.05 L(0.94) <0.05 L(0.69)
activemq-5.3.0 0.86 0.82 0.83 <0.05 L(0.85) <0.05 L(0.64)
activemq-5.8.0 0.88 0.85 0.86 <0.05 L(0.90) <0.05 L(0.80)
derby-10.2.1.6 0.88 0.74 0.85 <0.05 L(1.0) <0.05 L(0.88)
derby-10.3.1.4 0.88 0.84 0.85 <0.05 L(0.97) <0.05 L(0.90)
derby-10.5.1.1 0.87 0.82 0.84 <0.05 L(0.95) <0.05 L(0.77)
groovy-1.5.7 0.83 0.82 0.79 <0.05 S(0.17) <0.05 S(0.17)
groovy-1.6.BETA_1 0.90 0.82 0.87 <0.05 L(0.88) <0.05 L(0.54)
groovy-1.6.BETA_2 0.94 0.85 0.91 <0.05 L(0.90) <0.05 M(0.4)
hbase-0.94.0 0.86 0.72 0.84 <0.05 L(1.0) <0.05 M(0.45)
hbase-0.95.0 0.82 0.72 0.81 <0.05 L(1.0) <0.05 L(0.48)
hbase-0.95.2 0.73 0.71 0.71 <0.05 L(0.97) <0.05 L(0.54)
hive-0.10.0 0.86 0.72 0.83 <0.05 L(1.0) <0.05 L(0.75)
hive-0.12.0 0.86 0.80 0.84 <0.05 L(1.0) <0.05 L(0.54)
hive-0.9.0 0.91 0.79 0.9 <0.05 L(1.0) <0.05 L(0.53)
jruby-1.1 0.88 0.86 0.85 <0.05 S(0.25) <0.05 M(0.44)
jruby-1.4.0 0.86 0.83 0.84 <0.05 L(0.56) <0.05 L(0.49)
jruby-1.5.0 0.87 0.84 0.86 <0.05 L(0.59) <0.05 M(0.41)
jruby-1.7.0.preview1 0.86 0.84 0.83 <0.05 L(0.59) <0.05 L(0.74)
lucene-2.3.0 0.94 0.85 0.91 <0.05 L(1.0) <0.05 L(0.83)
lucene-2.9.0 0.84 0.79 0.82 <0.05 L(0.96) <0.05 L(0.66)
lucene-3.0.0 0.92 0.91 0.9 <0.05 S(0.32) <0.05 L(0.62)
lucene-3.1.0 0.79 0.74 0.76 <0.05 L(0.91) <0.05 M(0.44)
wicket-1.3.0-beta2 0.85 0.81 0.82 <0.05 L(0.76) <0.05 L(0.61)
wicket-1.3.0-incubating-beta-1 0.88 0.83 0.87 <0.05 L(0.82) <0.05 M(0.44)
wicket-1.5.3 0.82 0.81 0.79 <0.05 S(0.24) <0.05 L(0.65)
Median 0.87 0.82 0.84
[…]
Observation 9) Combining KU-based features with traditional code metrics results in a higher performing model for predicting post-release defects. Figure 9 depicts the performance of the studied models. We observe that the combined model, KUCLS+CC, is the top performing model, surpassing all the studied models including the individual KUCLS and CC for predicting post-release defects. The median AUC of the KUCLS+CC is 0.87, representing a normalized improvement of 16.8% over the CC, 27.8% over the KUCLS and 35% over the CC_PROD. We further examine the performance of KUCLS+CC for individual releases. Table 5 presents the P-value, effect size, and median AUC for the KUCLS+CC, CC and KUCLS models for each studied release. We observe that the KUCLS+CC models outperform the CC and the KUCLS in all 28 studied releases, with the majority showing a large effect size (highlighted in green in the columns “KUCLS+CC vs KUCLS” and “KUCLS+CC vs CC”). We also calculate the normalized AUC improvement for each of the studied releases. The normalized AUC improvement of the KUCLS+CC over the CC ranges from 4.9% to 33.3%, while the improvement over the KUCLS ranges from 5.6% to 59.9%. We thus conclude that the the KUCLS+CC demonstrates superior performance for predicting post-release defects, making it more effective than either individual models.
Figure 10: The rank distribution of features of the KUCLS+CC for defect classification. The number inside the square brackets indicates the final rank of the feature after applying Scott-Knott ESD for the second time. The KUs are highlighted with the blue colored font and traditional code metrics are highlighted with black colored font.
Observation 10) DDEV is the top ranked important feature for predicting post-release defects while Inheritance and Method Encapsulation of the KU-based features share the rank as the third most important features for the KUCLS+CC. Figure 10 presents the rank distribution of each feature (final rank is shown in brackets) of the KUCLS+CC across the studied releases. The top two important features are the process metrics: DDEV (i.e., the number of distinct developers) is ranked first, followed by Added Lines (the normalized count of lines added to the module) as the second most important feature. We then have three features that are ranked third most important features. Two of these features (Inheritance and Method Encapsulation) are the KU-based features and the other one is the product metric (i.e., ConntClassCoupled). The inclusion of Inheritance and Method Encapsulation KUs among the top three features highlights the significant contribution of KU-based features to the performance of the KUCLS+CC in predicting post-release defects.
Summary RQ4: Can KUCLS be made more accurate by combining KU-based features with traditional code metrics? Yes. We combined the features of KUCLS and CC models (KU-based features and traditional code metrics) and build a new combined model that we call KUCLS+CC. The KUCLS+CC outperforms the CC and KUCLS with a non-negligible effect size. In particular: KUCLS+CC ranks first among the studied models (CC, KUCLS and CC_PROD models) with a median AUC of 0.87. The normalized AUC improvement of the KUCLS+CC over the CC ranges from 4.9% to 33.3% and the improvement over the KUCLS ranges from 5.6% to 59.9% across the studied releases. In the KUCLS+CC, DDEV emerges as the most influential feature overall. Among the KU-based features, Inheritance and Method Encapsulation KUs are the third ranked features indicating significant contribution of KU-based feature in predicting post-release defects.
5.5 RQ5: How different classification algorithms with hyper-parameter tuning impact the combined model for predicting post-release defects?
5.5.1 Motivation
We use a random forest classifier with default hyper-parameter settings as the standard configuration for building the combined model KUCLS+CC (Section 5.4). However, the performance of a model can vary significantly across different classifiers, as each classifier possesses unique strengths and weaknesses tailored to specific data characteristics and problem types (Tantithamthavorn and Hassan, 2018). Moreover, achieving optimal performance often depends on fine-tuning a classifier’s hyper-parameters (Tantithamthavorn et al., 2016; Liao et al., 2022; Wong et al., 2019). Proper tuning can enhance the model’s capacity to learn from training data more effectively and efficiently. Therefore, in this research question, we explore whether a more accurate version of KUCLS+CC can be achieved by replacing the default random forest classifier with an alternative classifier that has been optimized through hyper-parameter tuning.
Table 6: The studied classification algorithms with different parameter settings for hyper parameter optimization.
Classifier
Name
Classification
Algorithm
Parameter
Name
Parameter
Description
Studied candidate
parameter values
KNN
K-Nearest
Neighbour
n_neighbors
The number of neighbors
required for each sample
{1, 5, 9, 13, 17, 20}
NB Naive Bayes var_smoothing
The portion of the largest variance
of all features that is added to
variances for calculation stability.
{1e-5, 1e-9, 1e-11, 1e-15}
DT Decision Tree criterion
The function to measure the quality
of a split.
{‘gini’, ‘entropy’, ‘log_loss’}
max_depth The maximum depth of the tree. {None, 5, 10}
ccp_alpha
Complexity parameter used for
Minimal Cost-Complexity Pruning.
{0.0001, 0.001, 0.01, 0.1, 0.5}
RF Random Forest n_estimator The number of trees in the forest. {10, 50, 100, 200}
max_depth
The maximum depth of the tree.
If None, then nodes are expanded
until all leaves are pure.
{None, 5, 10}
XGB XGBoost n_estimator
The number of boosting rounds or
trees to build.
{10, 50, 100, 200}
max_depth The maximum depth of the tree. {None, 5, 10}
learning_rate
This parameter controls how much
the model is adjusted in response
to the estimated error each time
the model weights are updated.
{0.1, 0.01, 0.001}
LGBM LightGBM n_estimator The number of boosted trees to fit. {10, 50, 100, 200}
num_leaves
The maximum tree leaves for
base learners.
{None, 5, 10}
learning_rate
The Boosting learning rate that
controls how quickly the model
adjusts to the error in each
iteration of training
{0.1, 0.01, 0.001}
5.5.2 Approach
To evaluate the impact of different classifiers and hyper-parameter tuning on the performance of KUCLS+CC, we experiment with a diverse set of classifiers: Naive Bayes (NB), K-Nearest Neighbor (KNN), Decision Trees (DT), Random Forest (RF), XGBoost (XGB), and LightGBM (LGBM). XGB and LGBM represent state-of-the-art classifiers known for their advanced capabilities, while the other classifiers are well-established and widely used in empirical software engineering research.
Figure 11: The AUC distribution of hyper-parameter-tuned models and our original KUCLS+CC (random forest with default hyper-parameter values). The names of hyper-parameter tuning models begin with “HPT_X”, where X represents the name of one of the classifier we studied. The models are grouped based on their performance rankings determined by the Scott-Knott ESD (SK-ESD) method, where a lower SK-ESD rank indicates a better-performing model.
We employ the out-of-sample bootstrap model validation technique with 100 repetitions to ensure robust evaluation. In each iteration, a bootstrap sample is generated to train the model, and the model is tested on the out-of-sample data (i.e., data not included in the bootstrap sample). To identify the optimal configuration for each classifier, we use Scikit-learn’s GridSearchCV [^12] https://scikit-learn.org/stable/modules/generated/sklearn.model_selection.GridSearchCV.html. This method performs an exhaustive search over a predefined grid of hyper-parameters (detailed in Table 6) and employs 10-fold cross-validation to evaluate each configuration. GridSearchCV is applied to the bootstrap sample data, identifying the model configuration that achieves the best performance based on hyper-parameter tuning. The best-tuned model is then tested on the out-of-sample data.
For model evaluation, we follow the methodology described in Step 1.4 of Section 5.1, ensuring consistency in assessing the effectiveness of the classifiers and their tuned configurations.
5.5.3 Findings
Observation 11) Our original KUCLS+CC models remain the top performing models for predicting post-release defects, outperforming all other hyper-parameter tuned classifiers. Figure 11 presents the AUC distribution of KUCLS+CC and the hyper-parameter-tuned variations. The KUCLS+CC model and the hyper-parameter-tuned random forest model (HPT_RF_KUCLS+CC) perform similarly, both leading in ranking. In contrast, models built using traditional classifiers such as Naive Bayes (NB), Decision Trees (DT), and K-Nearest Neighbor (KNN) exhibit suboptimal performance, with median AUC scores below 0.80. Of these, the KNN-based model performs the worst, with a median AUC below 0.76. Advanced classifiers like XGBoost and LightGBM achieve better results, with median AUC scores exceeding 0.85, yet they still fail to surpass the performance of the original KUCLS+CC model. Thus, simply switching to a different classifier with tuned hyper-parameters does not improve the performance of KUCLS+CC.
Summary RQ5: How different classification algorithms with hyper-parameter tuning impact the combined model for predicting post-release defects? We can not improve the performance of our original KULCS+CC models by simply using a different classification algorithm with tuned hyper-parameters.
5.6 RQ6: How does the performance of a cost-effective combined model compare to that of the model built with traditional code metrics?
5.6.1 Motivation
Building predictive models with a broad set of features often yields high performance, but this approach can be resource-intensive, requiring significant efforts in data collection and processing. These challenges highlight the importance of developing cost-effective models that achieve strong predictive performance while relying on a minimal, carefully selected set of features. Such models not only reduce resource requirements but also enhance practicality and scalability in real-world.
5.6.2 Approach
We focus on a carefully selected subset of KU-based features and traditional code metrics to build a cost-effective combined model. Specifically, we select the top five KU-based features (out of 28) from the KUCLS and the top five code metrics (out of 59) from the CC. This results in the construction of our cost-effective model, KUCLS_CC_COST_EFF, using only 10 features – a reduction of 89% in the total number of features. The selected features are as follows: Inheritance KU, Method Encapsulation KU, Exception KU, Advanced Class Design KU, Loop KU, CountClassCoupled, CountLineComment, MaxNesting_Mean, DDEV, and Added Lines.
To build the KUCLS_CC_COST_EFF models, we follow the approach for model construction presented in Section 5.1. We determine the best classifier (with hyper-parameter tuning) that performs the best in predicting LTCs by applying the same approach described in RQ4. Then, we evaluate KUCLS_CC_COST_EFF models and compare the performance of them with that of KUCLS, CC and KUCLS+CC models. The comparison involves ranking the models according to the same SK-ESD method detailed in Section 5.2.
Figure 12: The distribution of AUC of KUCLS_CC_COST_EFF, KUCLS, CC, and the KUCLS+CC across the studied releases. The models are grouped based on their performance rankings determined by the Scott-Knott ESD (SK-ESD) method, where a lower SK-ESD rank indicates a better-performing model.
5.6.3 Findings
Observation 12) The cost effective KUCLS_CC_COST_EFF model outperforms the CC model (analogous to KUCLS+CC) . Figure 12 presents the AUC distribution of the KUCLS_CC_COST_EFF, KUCLS, CC, and KUCLS+CC models. Upon rerunning the approach from RQ4, we observe that random forest with default parameter settings is the best classifier that we refer to the cost effective model as KUCLS_CC_COST_EFF. The KUCLS_CC_COST_EFF achieves a median AUC of 0.86, ranking second overall across the studied releases and consistently outperforming both the CC and KUCLS models.
We further analyze the model’s performance for each of the studied releases. Table 7 presents the P-value and effect size of the classification models for each studied release. The KUCLS_CC_COST_EFF outperforms the CC in most of the studied releases (25 out of 28), with a majority of these cases showing a large effect size. For the remaining releases, both models perform the same. Similarly, the KUCLS_CC_COST_EFF model outperforms the KUCLS model in most releases. The normalized AUC improvement of the KUCLS_CC_COST_EFF over the CC ranges from 5.5% to 33.3%, while the improvement over the KUCLS ranges from 7.7% to 61.9%. However, it is important to note that when budget constraints are not a consideration, the KUCLS+CC model remains the recommended choice, as it consistently ranks as the top-performing model (see the Rank-1 in Figure 12).
Table 7: P-value, effect size, and median AUC value of KUCLS_CC_COST_EFF, CC, KUCLS and KUCLS+CC defect classification models.
Statistical Test
KUCLS_CC_COST_EFF
vs CC
KUCLS_CC_COST_EFF
vs KUCLS
KUCLS_CC_COST_EFF
vs KUCLS+CC
Release P-value Effect Size P-value Effect Size P-value Effect Size
activemq-5.0.0 <0.05 L(0.72) <0.05 L(1.00) <0.05 S(-0.24)
activemq-5.1.0 <0.05 L(0.53) <0.05 M(0.41) >0.05 N(-0.05)
activemq-5.2.0 <0.05 L(0.93) <0.05 L(0.99) <0.05 L(0.59)
activemq-5.3.0 <0.05 L(0.71) <0.05 L(0.89) <0.05 N(0.05)
activemq-5.8.0 <0.05 M(0.42) <0.05 L(0.72) <0.05 L(-0.51)
derby-10.2.1.6 <0.05 L(0.79) <0.05 L(1.00) <0.05 S(-0.23)
derby-10.3.1.4 <0.05 L(0.58) >0.05 N(-0.074) <0.05 L(-0.99)
derby-10.5.1.1 <0.05 S(0.28) <0.05 L(0.77) <0.05 L(-0.61)
groovy-1.5.7 <0.05 N(-0.08) <0.05 S(-0.26) <0.05 S(-0.23)
groovy-1.6.BETA_1 <0.05 L(0.61) <0.05 L(0.91) <0.05 N(0.08)
groovy-1.6.BETA_2 >0.05 N(-0.01) <0.05 L(0.73) <0.05 M(-0.40)
hbase-0.94.0 <0.05 S(0.29) <0.05 L(1.00) <0.05 S(-0.20)
hbase-0.95.0 <0.05 L(0.94) <0.05 L(1.00) <0.05 L(0.70)
hbase-0.95.2 <0.05 L(0.72) <0.05 L(0.99) <0.05 S(0.26)
hive-0.10.0 <0.05 L(0.91) <0.05 L(1.00) <0.05 M(0.38)
hive-0.12.0 <0.05 L(0.81) <0.05 L(1.00) <0.05 M(0.40)
hive-0.9.0 <0.05 L(0.69) <0.05 L(1.00) <0.05 S(0.20)
jruby-1.1 <0.05 L(0.63) <0.05 M(0.45) <0.05 S(0.24)
jruby-1.4.0 <0.05 S(0.29) <0.05 M(0.40) <0.05 S(-0.24)
jruby-1.5.0 <0.05 M(0.42) <0.05 L(0.58) >0.05 N(0.05)
jruby-1.7.0.preview1 <0.05 L(0.87) <0.05 L(0.78) <0.05 S(0.31)
lucene-2.3.0 <0.05 L(0.88) <0.05 L(1.00) <0.05 N(0.08)
lucene-2.9.0 <0.05 M(0.36) <0.05 L(0.88) <0.05 M(-0.35)
lucene-3.0.0 <0.05 S(0.16) <0.05 S(-0.18) <0.05 L(-0.50)
lucene-3.1.0 <0.05 N(-0.09) <0.05 L(0.66) <0.05 L(-0.53)
wicket-1.3.0-beta2 <0.05 S(0.29) <0.05 L(0.50) <0.05 M(-0.36)
wicket-1.3.0-incubating-beta-1 <0.05 L(0.73) <0.05 L(0.92) <0.05 M(0.43)
wicket-1.5.3 <0.05 L(0.48) >0.05 N(0.02) <0.05 S(-0.24)
[…]
Figure 13: The rank distribution of features of KUCLS_CC_COST_EFF for defect classification. The number inside the square brackets indicates the final rank of the feature after applying Scott-Knott ESD for the second time. The KUs are highlighted with the blue colored font and traditional code metrics are highlighted with black colored font.
Observation 13) Method Encapsulation stands out as a key KU-based feature in the KUCLS_CC_COST_EFF model, ranking as the third most influential feature for predicting post-release defects. Figure 13 presents the rank distribution of each feature (final rank is shown in brackets) of the KUCLS_CC_COST_EFF across the studied releases. Similar to the feature importance results of the KUCLS+CC, the top two important features of the KUCLS_CC_COST_EFF are the code metrics (process metrics): DDEV and Added Lines. Among them, DDEV (number of distinct developers) is consistently ranked as the most important feature, with a median rank of 1, indicating its important role in predicting post-release defects. The KU-based feature Method Encapsulation is ranked third. Its narrow interquartile range, compared to most of the lower-ranked features, indicates consistent ranking across the studied releases. This result highlights the significant contribution of KU-based feature in the cost-effective model. Other notable features include CountLineComment which ranks fourth, while CountClassCoupled, and Exception Handling share the fifth rank. Additionally, features like Inheritance, Loop, and Advanced Class Design are ranked lower, but indicating significant contributions for predicting post-release defects. This analysis underscores the balanced integration of code metrics and KU-based features in achieving a cost-effective yet high-performing predictive model.
Summary RQ6: How does the performance of a cost-effective combined model compare to that of the model built with traditional code metrics? The cost-effective model KUCLS_CC_COST_EFF outperforms the CC (models built with traditional code metrics). In particular, The random classifier with default parameter settings achieve the best performance for KUCLS_CC_COST_EFF with a median AUC of 0.86. The normalized AUC improvement of the KUCLS_CC_COST_EFF over the CC ranges from 5.5% to 33.3%. Method Encapsulation is ranked as the third most influential feature, underscoring the significant contribution of KU-based features in the performance of the KUCLS_CC_COST_EFF (similar to the KUCLS).
6 Discussion
In this section, we discuss the improved predictive power of combining KUs with traditional metrics, examines the trade-offs between model performance and complexity, and outlines future directions for expanding the concept of KUs to include domain-specific and technical knowledge.
The complementary role of KUs in enhancing defect prediction with traditional code metrics. In our preliminary study, we observe that KUs exhibit minimal correlation with traditional code metrics, indicating that they capture different aspects of software systems. Additionally, the clusters of files generated using KUs do not overlap with those generated using code metrics, further emphasizing their distinctiveness. In RQ4, we find that combining KUs with traditional code metrics results in a higher-performing model (KUCLS+CC) that outperforms models built solely with traditional metrics, such as CC_PROD and CC. This demonstrates that integrating KUs provides more accurate identification of defect-prone code. Furthermore, our instance-specific analysis in RQ3 reveals that KUs offer complementary insights into defect-prone code that traditional metrics alone cannot provide. Specifically, KUs can capture traits of the software systems that are tied to the building blocks of the programming language (e.g., constructs and APIs) and provide insights of the defect-prone code related to these building blocks of the programming language. These findings highlight the importance of using KUs alongside traditional metrics to improve defect prediction.
Balancing model performance and complexity. We observe that the combined model, KUCLS+CC, outperforms all the studied defect prediction models. The normalized AUC improvement of the KUCLS+CC over the CC ranges from 4.9% to 33.3% and the improvement over the KUCLS ranges from 5.6% to 59.9% across the studied releases. However, this performance improvement comes with an increase in model complexity, particularly in terms of data processing and feature engineering. For example, KUCLS+CC combines 28 KU-based features with 59 traditional code metrics, resulting in a total of 87 features used to build the model, which can add significant overhead. To reduce such operational bur- dens, we introduce a cost-effective predictive model (KUCLS_CC_COST_EFF) that also outperforms the CC and KUCLS models. This cost-effective model only consists of 10 features (five from the KU-based and five from the traditional code metrics), makes it a far more cheaper than KUCLS+CC. We want to note that when cost constraints are not a concern, the KUCLS+CC should be the recommended model. It not only effectively identifies defects but also provides insights into defects related to all categories of KUsKUs, including rarely used (e.g., Message Service KU) and advanced KUs (e.g., Concurrency and IO KUs).
Expanding KUs of the programming language to include other forms of knowledge (i.e., domain knowledge and library knowledge). In this paper we leverage KUs of the Java programming language to predict post-release defects. Models built based on the KU-based features (e.g., KUCLS) and the combination of them with traditional code metrics (e.g., KUCLS+CC) outperform the models built with traditional code metrics. Such results indicate that KUs are good predictors for defect prediction.
Given our encouraging results in defect prediction, researchers should explore the KUs of other programming languages (e.g., python and Ruby). Fortunately, certification exams exist for several languages and should thus serve as a useful starting point. Additionally, researchers should consider expanding KUs to include other forms of knowledge, such as domain-specific knowledge units. A promising approach could involve analyzing project descriptions from README files and developer discussions in pull requests or issue trackers, leveraging large language models (LLMs) to summarize and extract relevant domain knowledge. We also encourage future works to integrate other form of technical knowledge such as libraries and API knowledge. This could include high-level libraries widely used in data science, such as pandas, as well as lower-level, such as OpenSSL for cryptographic operations, and libcurl for network communication. Incorporating these diverse forms of knowledge could provide deeper insights into code behavior and can significantly enhance the predictive power of models for identifying post-release defects.
7 Threats to Validity
Construct validity. In Section  4, we applied PCA on the KU-based feature data and code metrics data to reduce the dimensionality of the data before running the k-means algorithm. To reduce the dimensionality of the data, one needs to select few PCA components. However, selecting the right number of PCA components is crucial for the study as selecting a large number of PCA components can still retain the same dimensionality problem. Selecting a very small number of PCA components, albeit reducing the dimensions of the data, can hinder the explainability of the data. To select these PCA components, we decided to use a threshold-based approach on the explained variance of the PCA components. We selected the first few PCA components that aggregate up to the 95% of the data which is a reasonable threshold for the explained variance of the data (Xibilia et al., 2020; Ghotra et al., 2017; Kondo et al., 2019).
We use the Java certification exams as guidelines to identify KUs because certification exams cover important topics of the Java programming language. We only focus on the core functionalities of the programming language that are listed in the topics of these exams. The set of KUs that we conceive in this paper is thus limited and does not cover all the knowledge embedded in the constructs and APIs of the Java programming language. However, these core functionalities are covered in the Java certification exams (Java SE and Java EE) to assess the expertise of developers. We encourage future studies to refine our conceptualization of KUs and operationalize them in different ways.
Another construct validity threat stems from the fact that we did not extract KUs from the third-party libraries used by the studied systems. In this study, we only focus on those KUs that are implemented by the developers of the studied software systems. We carefully implemented the type binding resolutions using the Java JDT (Eclipse, 2020) to ensure that we are less likely to miss the KUs implemented within the studied projects.
Internal validity. The results reported in this paper are based on models that were trained without applying data rebalancing techniques, such as SMOTE (Chawla et al., 2002). To investigate whether class rebalancing techniques could improve model performance, we applied SMOTE and re-ran our models. However, we observed no improvement in performance. Moreover, such class rebalancing techniques are not recommended when building explainable models, as they can introduce issues like concept drift and hinder model interpretability (Tantithamthavorn and Hassan, 2018). Therefore, we chose to refrain from using SMOTE in our study.
External validity. In this study, we only considered Java projects. Future studies should broaden the scope of our study and investigate how our findings apply to software projects written in other popular programming languages that have certification exams in different development areas (e.g., web development). For instance, Ruby on Rails, JavaScript, Python, ASP.Net, Node-js, and Objective-C are the top languages for web developers and all of these languages have their own certification exams (Vekariya, 2018).
8 Related Work
A number of studies have focused on analyzing different metrics for defect prediction models (Briand et al., 2000; Emam et al., 2001; Zhou and Leung, 2006; Gyimothy et al., 2005; Nagappan et al., 2006, 2006; Zhang, 2009; Olague et al., 2007; Kanmani et al., 2007; Moser et al., 2008; Aggarwal et al., 2009; Burrows et al., 2010; Kumar et al., 2017). In this section, we highlight a few of these studies that are more closely related to our study.
The simplest, easiest to extract, and most commonly used product metric for bug prediction is the lines of code (LOC). Zhang (2009) investigated the relationship between LOC and bug in Eclipse and several NASA projects. The author observed that typical classification techniques based on LOC were able to predict defective components reasonably well. Koru et al. (2008, 2008) observed a power-law relationship between LOC and bug-proneness, with the latter increasing at a slower rate. They also observed that smaller modules were proportionally more bug-prone compared with larger ones. Although LOC correlates with the number of bugs (Ostrand et al., 2005; Zhang, 2009; English et al., 2009), some studies find no strong evidence that LOC is a good indicator of bugs (Fenton and Ohlsson, 2000; Andersson and Runeson, 2007). Many researchers investigated other product metrics based on code, such as object-oriented metrics and complexity metrics for predicting bugs. These metrics are commonly referred to as code metrics. Gyimothy et al. (2005) used eight object-oriented metrics and traditional code-size metrics (e.g., LOC) to detect bug-proneness in the Mozilla source code. Their results indicated that all code metrics (except the one that calculates the number of direct descendants of a class) were significant for bug prediction. Nagappan et al. (2006) built prediction models using 18 complexity metrics along three dimensions (module, function, and class). Their evaluation results showed that complexity metrics were able to predict post-release defects with good accuracy, but no single set of metrics was applicable to all projects.
Many prior studies build bug-prediction models with process metrics and investigate the performance of the bug prediction task between models that are built with product metrics and process metrics. Nagappan et al. (2006) investigated the utility of using process and product metrics to estimate post-release bug-proneness in Windows XP-SP1 and Windows 2003 Server operating systems. They observed that process and product metrics could be used to identify bug-prone modules at statistically significant levels. Moser et al. (2008) proposed 18 process metrics based on the change history of the files (i.e., change metrics). They studied the performance of bug prediction models built with these 18 process metrics and 31 traditional static code metrics. They observed that predictors based on process metrics outperformed predictors based on static code metrics. In contrast with the file-level bug prediction approach, Giger et al. (2012) explored bug prediction models at the method level. They developed prediction models using a combination of 15 process metrics and 9 source code metrics to assess their performance on 21 Java projects. Their findings indicated that process metrics alone were effective predictors and significantly outperformed source code metrics. However, combining both types of metrics did not enhance the models’ classification performance. Rahman and Devanbu (2013) built defect prediction models using 54 code metrics and 14 process metrics and evaluated their performance on 85 releases of 12 major open-source projects. The results of their study revealed that code metrics were generally less effective than process metrics.
Majumder et al. (2022) carried out an extensive study analyzing 722K commits from 700 GitHub projects to compare the effectiveness of bug prediction models that use process metrics against those that use product metrics. Their findings indicated that process metrics are superior in predicting defects. However, Dalla Palma et al. (2021) observed that product metrics are better than process metrics for identifying bugs in Infrastructure-as-Code (IaC) scripts, which are used in DevOps to manage and provision infrastructure through machine-readable files.
In this paper, we empirically study knowledge units (KUs) of programming languages, a new perspective for analyzing the source code of software systems. Our goal is to understand if we can leverage KUs for predicting post-release defects. Our findings indicate that KUs are good candidates for identifying defect-prone code and offer insights into the occurrence of defects or the context in which those defects happen.
9 Conclusion
In this paper, we present an empirical study on how to leverage KUs of a programming language to predict post-release defects. Our results show that KUs differ from traditional code metrics and that they can offer a new lens through which the code of software systems can be analyzed. With this encouraging results, we develop a defect prediction model called KUCLS, which uses KU-based features. We compare the performance of the KUCLS with the models built with traditional code metrics: CC_PROD (models built with product metrics) and CC (models built with product and process metrics). The KUCLS significantly outperforms the CC_PROD with a normalized AUC improvement ranges from 5.1% to 28.9%. Combining KUs and traditional code metrics lead to even a higher performing model (KUCLS+CC) that significantly outperforms the CC and KUCLS models. The normalized AUC improvement of the KUCLS+CC over the CC ranges from 4.9% to 33.3% and the improvement over the KUCLS ranges from 5.6% to 59.9% across the studied releases. We also observe that KUs provide insights into the incidence of post-release defects that are complementary to those offered by the traditional code metrics. Our encouraging results highlights the importance of KU-based features in defect prediction. Given our promising findings in this study, we encourage researchers to further explore and expand the concept of KUs. This includes (i) developing more advanced conceptualizations and operationalizations of KUs, such as those supporting the detection of various design patterns, and (ii) evaluating whether the insights from our study can be generalized to other software systems and programming languages.
Data Availability Statement (DAS)
A supplementary material package is provided online in the following link: http://bit.ly/3GoSmHL. The contents will be made available on a public GitHub repository once the paper is accepted.
APPENDICES
Appendix A Java Certification Exams and Knowledge Units
Table 8 summarizes our KUs and their capabilities.
Table 8: Knowledge Units derived from the Java SE 8 Programmer I Exam, Java SE 8 Programmer II Exam, and Java EE Developer Exam
Knowledge Unit (KU) Key Capabilities
[K1] Data Type [C1] Declare and initialize different types of variableS(e.g., primitive type, parameterized type, and array type), including the casting of primitive data types
[K2] Operator and Decision [C1] Use Java operators(e.g., assignment and postfix operators); use parentheses to override operator precedence
[C2] Test equality between strings and other objects using == and equals()
[C3] Create and use if, if-else, and ternary constructs
[C4] Use a switch statement
[K3] Array [C1] Declare, instantiate, initialize and use a one-dimensional array
[C2] Declare, instantiate, initialize and use a multi-dimensional array
[K4] Loop [C1] Create and use while loops
[C2] Create and use for loops, including the enhanced for loop
[C3] Create and use do-while loops
[C4] Use break statement
[C5] Use continue statement
[K5] Method and Encapsulation [C1] Create methods with arguments and return values
[C2] Apply the “static” keyword to methods, fields, and blocks
[C3] Create an overloaded method and overloaded constructor
[C4] Create a constructor chaining (use “this()” method to call one constructor from another constructor
[C5] Use variable length arguments in the methods
[C6] Use different access modifiers (e.g., private and protected) other than “default”
[C7] Apply encapsulation: identify set and get method to initialize any private class variables
[C8] Apply encapsulation: Immutable class generation-final class and initialize private variables through the constructor
[K6] Inheritance [C1] Use basic polymorphism (e.g., a superclass refers to a subclass)
[C2] Use polymorphic parameter (e.g., pass instances of a subclass or interface to a method)
[C3] Create overridden methods
[C4] Create “abstract” classes and “abstract” methods
[C5] Create “interface” and implement the interface
[C6] Use “super()” and the “super” keyword to access the members(e.g., fields and methods) of a parent class
[C7] Use casting in referring a subclass object to a superclass object
[K7] Advanced Class Design [C1] Create inner classes, including static inner classes, local classes, nested classes, and anonymous inner classes
[C2] Develop code that uses the final
[C3] Use enumerated types including methods and constructors in an “enum” type
[C4] Create singleton classes and immutable classes
[K8] Generics and Collection [C1] Create and use a generic class
[C2] Create and use ArrayList, TreeSet, TreeMap, and ArrayDeque
[C3] Use java.util.Comparator and java.lang.Comparable interfaces
[C4] Iterate using forEach methods of List
[K9] Functional Interface [C1] Use the built-in interfaces included in the java.util.function packages such as Predicate, Consumer, Function, and Supplier
[C2] Develop code that uses primitive versions of functional interfaces
[C3] Develop code that uses binary versions of functional interfaces
[C4] Develop code that uses the UnaryOperator interface
[K10] Stream API [C1] Develop code to extract data from an object using peek() and map() methods, including primitive versions of the map() method
[C2] Search for data by using search methods of the Stream classes, including findFirst, findAny, anyMatch, allMatch, noneMatch
[C3] Develop code that uses the Optional class
[C4] Develop code that uses Stream data methods and calculation methods
[C5] Sort a collection using Stream API
[C6] Save results to a collection using the collect method
[C7] UseflatMap() methods in the Stream API
Knowledge Unit (KU) Key Capabilities
[K11] Exception [C1] Create a try-catch block
[C2] Use catch, multi-catch, and finally clauses
[C3] Use autoclose resources with a try-with-resources statement
[C4] Create custom exceptions and autocloseable resources
[C5] Create and invoke a method that throws an exception
[C6] Use common exception classes and categories(such as NullPointerException, ArithmeticException, ArrayIndexOutOfBoundsException, ClassCastException)
[C5] Use assertions
[K12] Date Time API [C1] Create and manage date-based and time-based events including a combination of date and time into a single object using LocalDate, LocalTime, LocalDateTime, Instant, Period, and Duration
[C2] Formatting date and times values for using different timezones
[C3] Create and manage date-based and time-based events using Instant, Period, Duration, and Temporal Unit
[C4] Create and manipulate calendar data using classes from java.time.LocalDateTime, java.time.LocalDate, java.time.LocalTime, java.time.format.DateTimeFormatter, and java.time.Period
[K13] IO [C1] Read and write data using the console
[C2] Use BufferedReader, BufferedWriter, File, FileReader, FileWriter, FileInputStream, FileOutputStream, ObjectOutputStream, ObjectInputStream, and PrintWriter in the java.io package
[K14] NIO [C1] Use the Path interface to operate on file and directory paths
[C2] Use the Files class to check, read, delete, copy, move, and manage metadata a file or directory
[K15] String Processing [C1] Search, parse and build strings
[C2] Manipulate data using the StringBuilder class and its methods
[C3] Use regular expression using the Pattern and Matcher class
[C4] Use string formatting
[K16] Concurrency [C1] Create worker threads using Runnable, Callable and use an ExecutorService to concurrently execute tasks [C2] Use synchronized keyword and java.util.concurrent.atomic package to control the order of thread execution
[C3] Use java.util.concurrent collections and classes including CyclicBarrier and CopyOnWriteArrayList
[C4] Use parallel Fork/Join Framework
[K17] Database [C1] Describe the interfaces that make up the core of the JDBC API, including the Driver, Connection, Statement, and ResultSet interfaces
[C2] Submit queries and read results from the database, including creating statements, returning result sets, iterating through the results, and properly closing result sets, statements, and connections
[K18] Localization [C1] Read and set the locale by using the Locale object
[C2] Build a resource bundle for each locale and load a resource bundle in an application
[K19] Java Persistence [C1] Create JPA Entity and Object-Relational Mappings (ORM)
[C2] Use Entity Manager to perform database operations, transactions, and locking with JPA entities
[C3] Create and execute JPQL statements
[K20] Enterprise Java Bean [S1] Create session EJB components containing synchronous and asynchronous business methods, manage the life cycle container callbacks, and use interceptors.
[S2] Create EJB timers
[K21] Java Message Service API [S1] Implement Java EE message producers and consumers, including Message-Driven beans
[S2] Use transactions with JMS API
[K22] SOAP Web Service [S1] Create SOAP Web Services and Clients using JAX-WS API
[S2] Create marshall and unmarshall Java Objects by using JAXB API
[K23] Servlet [S1] Create Java Servlet and use HTTP methods
[S2] Handle HTTP headers, parameters, cookies
[S3] Manage servlet life cycle with container callback methods and WebFilters
[K24] Java REST API [S1] Apply REST service conventions
[S2] Create REST Services and clients using JAX-RS API
[K25] Websocket [S1] Create WebSocket Server and Client Endpoint Handlers
[S3] Produce and consume, encode and decode WebSocket messages
[K26] Java Server Faces [S1] Use JSF syntax and use JSF Tag Libraries
[S2] Handle localization and produce messages
[S3] Use Expression Language (EL) and interact with CDI beans
[K27] Contexts and Dependency Injection (CDI) [S1] Create CDI Bean Qualifiers, Producers, Disposers, Interceptors, Events, and Stereotypes
[K28] Batch Processing [S1] Implement batch jobs using JSR 352 API
Appendix B Traditional code metrics
The list of traditional code metrics employed in this study can be seen in Table 9. This list corresponds to all code metrics that the Understand tool can calculate. Table 10 presents the list of process metrics that we select for this study.
Table 9: The list of traditional code metrics that we select for this study.
Granularity Metric Name Understand (API) name Definition
File Average Cyclomatic Complexity AvgCyclomatic Average cyclomatic complexity for all nested functions or methods.
LeveL(37) Average Modified Cyclomatic Complexity AvgCyclomaticModified Average modified cyclomatic complexity for all nested functions or methods.
Average Strict Cyclomatic Complexity AvgCyclomaticStrict Average strict cyclomatic complexity for all nested functions or methods.
Average Essential Cyclomatic Complexity AvgEssential Average Essential complexity for all nested functions or methods.
Average Number of Lines AvgLine Average number of lines for all nested functions or methods.
Average Number of Blank Lines AvgLineBlank Average number of blanks for all nested functions or methods.
Average Number of Lines of Code AvgLineCode Average number of lines containing source code for all nested functions or methods.
Average Number of Lines with Comments AvgLineComment Average number of lines containing comments for all nested functions or methods.
Classes CountDeclClass Number of classes.
Class Methods CountDeclClassMethod Number of class methods.
Class Variables CountDeclClassVariable Number of class variables.
Function CountDeclFunction Number of functions.
Instance MethodS(NIM) CountDeclInstanceMethod Number of instance methods.
Instance VariableS(NIV) CountDeclInstanceVariable Number of instance variables.
Local Methods CountDeclMethod Number of local methods.
Local Default Visibility Methods CountDeclMethodDefault Number of local default methods.
Private MethodS(NPM) CountDeclMethodPrivate Number of local private methods.
Protected Methods CountDeclMethodProtected Number of local protected methods.
Public Methods CountDeclMethodPublic Number of local public methods.
Physical LineS(NL) CountLine Number of all lines.
Blank Lines of Code (BLOC) CountLineBlank Number of blank lines.
Source Lines of Code CountLineCode Number of lines containing source code. [aka LOC]
Declarative Lines of Code CountLineCodeDecl Number of lines containing declarative source code.
Executable Lines of Code CountLineCodeExe Number of lines containing executable source code.
Lines with Comments CountLineComment Number of lines containing comments.
Semicolons CountSemicolon Number of semicolons.
Statements CountStmt Number of statements.
Declarative Statements CountStmtDecl Number of declarative statements.
Executable Statements CountStmtExe Number of executable statements.
Max Cyclomatic Complexity MaxCyclomatic Maximum cyclomatic complexity of all nested functions or methods.
Max Modified Cyclomatic Complexity MaxCyclomaticModified Maximum modified cyclomatic complexity of nested functions or methods.
Max Strict Cyclomatic Complexity MaxCyclomaticStrict Maximum strict cyclomatic complexity of nested functions or methods.
Comment to Code Ratio RatioCommentToCode Ratio of comment lines to code lines.
Sum Cyclomatic Complexity (WMC) SumCyclomatic Sum of cyclomatic complexity of all nested functions or methods.
Sum Modified Cyclomatic Complexity SumCyclomaticModified Sum of modified cyclomatic complexity of all nested functions or methods.
Sum Strict Cyclomatic Complexity SumCyclomaticStrict Sum of strict cyclomatic complexity of all nested functions or methods.
Sum Essential Complexity SumEssential Sum of essential complexity of all nested functions or methods
Class Coupling Between ObjectS(CBO) CountClassCoupled Number of other classes coupled to.
LeveL(5) Number of ChildreN(NOC) CountClassDerived Number of immediate subclasses.
Depth of Inheritance Tree (DIT) MaxInheritanceTree Maximum depth of class in inheritance tree.
Lack of Cohesion in MethodS(LCOM) PercentLackOfCohesion Lack of Cohesion in Methods
Response for clasS(RFC) CountDeclMethodAll Number of methods, including inherited ones.
Method InputS(FANINN) CountInput {Min, Median, Max} Number of calling subprograms plus global variables read.
LeveL(12) OutputS(FANOUT) CountOutput {Min, Median, Max} Number of called subprograms plus global variables set.
PathS(NPATH) CountPath {Min, Median, Max} Number of possible paths, not counting abnormal exits or gotos.
Nesting MaxNesting {Min, Median, Max} Maximum nesting level of control constructs.
Table 10: The list of process metrics that we select for this study.
Metric Name Description
COMM The number of Git commits
ADDED_LINES The normalized number of lines added to the module
DEL_LINES The normalized number of lines deleted from the module
ADEV The number of active developers
DDEV The number of distinct developers